OpenStudy (anonymous):

factor the trinomial. 18c2 +12c -6

OpenStudy (mathmale):

\[18c^2 +12c -6 \] has three terms (being a trinomial!), and those three terms share a common factor. What is that common factor? Factor it out, please.

OpenStudy (anonymous):

is the common factor 3?

OpenStudy (mathmale):

Certainly 3 is part of the common factor, but is there a larger common factor ("greatest common factor")?

OpenStudy (mathmale):

Well, let's go ahead and factor out the 3:

OpenStudy (anonymous):

y = 18x^2 + 12x - 6 y = 6(3x^2 + 2x - 1) = 6(x + 1) (3x - 1)

OpenStudy (mathmale):

@kaayllllaa: Your thoughts? Please respond in some way.